### v2

# Topologically-Protected Remanent Vortices in Confined Superfluid 3He

Abstract
Symmetry-breaking phase transitions from less to more ordered phases will typically produce topological defects in the ordered phase. Kibble-Zurek theory predicts that for any second-order phase transition, such as the early universe, the density of defects that form should be determined by the scaling law for the system coherence time and the phase transition quench time. We have performed measurements of fourth sound dissipation due to vortex mutual friction in thin channels of superfluid 3He where one spatial dimension is smaller than a characteristic length scale predicted by the Kibble-Zurek theory. Our measurements suggest that remanent vortices form after the normal to superfluid second-order phase transition, and that the density of vortices is correlated with the size of the channel, but crucially, is independent of quench time. We propose a modified picture of defect formation, where closely spaced walls prevent the ends of vortex lines from reconnecting into loops. This leads to a mean vortex separation set by the wall spacing, which can result in much higher defect densities than in bulk systems.
preprint: APS/123-QED
I Introduction
Phase transitions, the processes by which the microscopic order of a system spontaneously changes, are of fundamental importance to many scientific fields, not only in chemistry and condensed matter physics [1] where the concept finds its origins, but also in cosmology [2], cell biology [3], and economics [4]. In thermodynamic terms, phase transitions can be broadly classified into first- and second-order phase transitions. First-order phase transitions are characterized by a discontinuity in the entropy, whereas second-order transitions feature a continuous, but not smooth, change in entropy which creates a discontinuity in the specific heat. Alternatively one may understand phase transitions in terms of the order parameter, which quantifies the degree to which the system exhibits long-range correlations in a physical quantity specific to the system (e.g., density, magnetic moment, etc.) [5]. Such classification naturally gives rise to a length scale known as the correlation length, over which the system is ordered. A general feature of second-order phase transitions is the divergence of the system’s correlation length at a critical temperature, representing the change from a disordered system (correlated only at very small distances) to one with long-range order.
Correlation cannot spread throughout a system instantaneously. Shortly after the system crosses the critical temperature, the equilibrium correlation length is limited by the speed at which information propagates through the system. In a discussion of phase transitions in the early universe, Kibble [6] noted that correlation in a field, such as the scalar Higgs field, must be limited by causality since information cannot propagate faster than the speed of light. This implied the existence of cosmic domains where field correlations may only exist within the domain. Because each domain develops independently from every other, the expectation value of the field in each domain will be independent. Correlation in the Higgs field spreads as these causally disconnected domains propagate outward at the speed of light. As the field evolves in time, the topology of the initial domain structure may be such that topological defects, such as domain walls, cosmic strings, or monopoles persist.
This argument was extended by Zurek, who discussed a similar process in the context of the phase transition from liquid to superfluid 4He. In this transition, the condensate wave function (i.e., the wavefunction of 4He atoms occupying the quantum ground state) becomes correlated at long distances [7]. Zurek took a new approach to estimating the density of defects by comparing the power-law divergence of the coherence length, $\xi$, and time, $\tau$, to the rate at which the system temperature is swept through the critical temperature. Here, the speed at which domains spread, $c$, is given by the speed at which fluctuations propagate. If the system temperature or pressure is swept through a critical value quickly, then these domains will have insufficient time to propagate, leading to the creation of string-like defects known as quantized vortices. This is distinct from the adiabatic limit, where the system is always at equilibrium, resulting in no vortices being nucleated by the phase transition.
The dynamics of phase transitions can be greatly simplified by assuming that changes in the order parameter occur sufficiently slowly that the system always remains near equilibrium. Non-adiabatic dynamics come into play when the transition between phases occurs quickly, relative to the rate at which information can propagate through the system. Quenching of the order parameter results in the freezing-in of numerous domains, the vertices of which represent string-like topological defects (vortices) in the ordered phase (the superfluid state). Kibble-Zurek theory describes this process in terms of an equilibrium correlation length $\xi$ and an equilibrium relaxation time $\tau$. If the quench is fast compared to the relaxation time of the system, then there will be insufficient time for the system to settle into a single translationally invariant domain. Instead, many domains will nucleate throughout the sample and grow, propagating outwards at a velocity characteristic to the system (e.g., the speed of entropy propagation in superfluid helium). The number of distinct domains that exist after the transition will depend on the amount of time these domains have to propagate. The number of domains that will form can be predicted by a comparison of the temperature ramp rate (or a rate in an analogous quantity that parameterizes the phase transition) to the relaxation time. Owing to the fact that the relaxation time diverges near the critical temperature, the domain growth is said to be effectively stopped after a “freeze-out” time $\hat{t}$. The original prediction of Kibble-Zurek theory is that the average size of a domain, $\hat{\xi}$, will be the size of the correlation length evaluated at the freeze-out time. A more technical description can be found in the Supplementary Material [8].
Figure 1: Domain Illustration. (a) During a rapid quench, domains with average size $\hat{\xi}$ form, giving rise to defects. Each domain will independently select a random phase $\phi$. If the phases $\phi_{1}$, $\phi_{2}$, and $\phi_{3}$ increase sequentially (either clockwise or counter-clockwise) then the resulting phase winding at the vertex will form a vortex. The density of vortices predicted to form in this area is of order $\hat{\xi}^{-2}$. (b) The average domain size in 3He for a temperature ramp rate of 0.12 mK/hour is predicted by Kibble Zurek theory to vary from 60-160 $\mu$m as a function of pressure [8]
Each domain in the ordered phase has a different microscopic structure, described by its order parameter, which randomly sets in upon formation of the domain. In general, the form of the order parameter is dependent on the symmetries of system in question. It may be a complex scalar quantity (e.g., in superfluid 4He), a vector (e.g., spin or liquid crystal systems), or can have a more complex tensor form (e.g., superfluid 3He). In all cases, the order parameter is continuous within a domain, but with singular points along the boundaries between domains. Such discontinuities in the order parameter field are called defects. A defect is said to be topological if removing it requires not merely local perturbation, but complete destruction of the macroscopic order [9]. In 3D systems, topological defects typically take the form of lines [9], and arise during a quench if the intersection of three or more domains has the correct topological structure. In the case of superfluid helium (both 3He and 4He), domains feature distinct phases of the condensate wavefunction. Suppose the phases increase sequentially (either clockwise or counter-clockwise) as one traces out a circle around the vertex of three or more domains. In that case, the vertex will become a vortex once the system reaches equilibrium (see Figure 1). The areal density of vortices generated by the Kibble-Zurek mechanism therefore is estimated to be $L=\hat{\xi}^{-2}$, based on the number of domains that are expected to form across a given surface area. The density of defects formed by phase transitions have been studied in a number of systems such as liquid crystals [10], Bose-Einstein condensates [12, 13, 11], strongly interacting Fermi gasses [14], colloidal monolayers [15], and quantum annealing devices [16].
Surprisingly, despite superfluid 4He being the original model system for Zurek’s theory, experiments in 4He have not been compatible with the theory [17, 18]. The observed density of vortices was at least two orders of magnitude smaller than predicted [18]. The reason for this discrepancy is currently not well understood. More success has been found in studying superfluid 3He. The order parameter of 3He has a more complex form, including not just a complex phase, but a matrix structure related to the angular momentum degrees of freedom of Cooper pairs [19]. This may result in multiple types of vortices [20], as well as domain wall structures [21, 22, 23]. The types of defects that form depend on the superfluid phase in question (i.e., 3He-A, 3He-B, etc.), as well as the boundary conditions of the system. Further discussion can be found in the Supplementary Material [8]. Bulk 3He quenching can be achieved by using neutron irradiation to locally heat the helium, temporarily entering the normal state before rapidly cooling through $T_{c}$ again. These experiments create short-lived networks of vortex lines, which have been detected via nuclear magnetic resonance [24, 25]. More recently, the formation of single and half-quantum vortices due to the KZM in the polar phase of 3He submerged in aerogel has been studied [26, 27, 28].
Figure 2: Helmholtz Resonator Data. (a) Resonance curves of the three different devices, with varying nanoscale confinements, measured during a temperature sweep carried out at 15 bar. Inset shows the Lorentizan resonance shape — after background subtraction — at a single temperature. From the center frequency, the superfluid density can be extracted [29]. (b) Additionally, the linewidth of each device can be extracted, plotted here as a function temperature. (c) Resulting vortex densities calibrated from the Helmholtz resonator linewidths and the mutual friction, see main text. Note the lack of temperature dependence, and the dependence on confinement. (d) Vortex lines form at the intersection of Kibble-Zurek domains during phase transition. If vortex lines bind to the channel walls they are topologically protected from decay, extending the lifetime of typically short-lived vortices under nanoscale confinement. Greater confinement results in a larger areal density, $L$, of topologically protected vortices.
Methods
We have designed an experiment sensitive to vortex dissipation that is distinct from previous studies in helium. It consists of nanofluidic devices, known as Helmholtz resonators, which have been described in detail elsewhere [30, 31, 29]. These devices allow for confinement of the helium in channels of thickness $H_{1}=636$ nm, $H_{2}=805$ nm, and $H_{3}=1067$ nm, respectively. Additional follow-up experiments were later performed with a $750$ nm device. Parallel-plate capacitors patterned onto the devices allow for an electrostatic force, which slightly compresses the center of the device, modifying the volume, and creating a pressure gradient down the channel. This drives superfluid motion in the channels, but notably does not couple to normal-fluid motion due to the viscous penetration depth being large compared to the thickness of the channel [8]. By driving this superfluid Helmholtz mode, we are able to extract two relevant parameters: a center frequency of the Helmholtz mode, $f_{0}$, the square of which is proportional to the superfluid density, $\rho_{s}\propto f_{0}^{2}$; and a linewidth, $\Delta f$, which characterizes the resonator’s dissipation. We observe phase transitions as kinks in the Helmholtz resonance frequency  [29].
We find that the devices exhibit linewidths of several Hz (see Figure 2), corresponding to quality factors of a few hundred depending on the temperature and geometry in question. The quality factor is many orders of magnitude lower than that measured in similar devices in 4He [30]. For this reason, the dissipation must be limited by a process intrinsic to the 3He system, rather than the quartz substrate (which should be much lower in the millikelvin temperature regime compared to the 1 K regime 4He experiment). We have identified a small number of possible dissipation mechanisms which are intrinsic to the superfluid phase of 3He: pair-breaking, vortex mutual friction, normal fluid slip, and second viscosity. Additionally, there are phase-specific effects associated with 3He textures. In prior publications [32] we have investigated the A-phase’s critical velocity by studying the resonance amplitude’s drive dependence. A linear regime exists where the amplitude is proportional to the drive, and a change in the slope of the amplitude-force curve marks a non-linear regime. All of the measurements in this publication were performed in the linear regime for both the A and B-phases. In the linear regime, we expect the superflow to be robust against pair-breaking. Calculations regarding normal fluid slip, second viscosity, and textural effects have been presented in the Supplementary Material [8]. We find that these effects are too small to explain the observed linewidth. Furthermore, these effects do not exhibit the observed temperature dependence of the linewidth. We assert that the best explanation for the observed dissipation in the linear regime is vortex mutual friction. The basis for this is that the temperature dependence of the ratio, $\Delta f/f_{0}^{2}$, is well fit by a function linear in the quantity $\alpha(T)$
$\frac{\Delta f(T)}{f_{0}^{2}(T)}=C_{1}\alpha(T)+C_{2}.$ (1)
Here, $\alpha$ is the mutual friction parameter associated with dissipative vortex motion, which is calculated following Ref. [33] using the theoretical model developed by [34, 35]
$\alpha(T)=\left(\frac{\rho_{s}}{\rho}\right)\left(\frac{\Delta}{k_{B}T}\right)^{-2}\frac{e^{-\Delta/k_{B}T}}{{a_{K}}}\coth\left(\frac{\Delta}{2k_{B}T}\right).$ (2)
Here, $\Delta(T)$ is the superfluid gap of the isotropic B-phase, and $a_{K}$ is a fitting factor taken from [33]. The mutual friction parameter $\alpha(T)$ diverges as $T\to T_{c}$, as can be seen in Figure 3.
When the experimentally measured ratio $\Delta f(T)/f_{0}^{2}(T)$ is plotted against the theoretical $\alpha(T)$, a clear linear trend is observed. To explain this trend we introduce a mutual friction model for our device outlined in the Supplementary Material [8]. This model predicts the constant prefactor should be given by
$C_{1}=2\pi\kappa L\left(\frac{A^{2}\ell\rho}{2a}\right)\left(\frac{1+\Sigma}{k_{p}}\right)=\gamma L,$ (3)
where $\kappa$ is the quanta of circulation, $A$ is the area of the Helmholtz resonator basin, $a$ is the cross-sectional area of the channel, $\ell$ is the length of the channel, $\rho$ is the total mass density of the fluid, $k_{p}$ is the spring constant of the quartz stiffness, and $\Sigma$ is the ratio of the quartz spring constant to the effective spring constant due to the helium compressabillity. From either of the fitting parameters, we can then extract the vortex density $L$
$L=\frac{C_{1}}{\gamma}=\frac{(\Delta f/f_{0}^{2}-C_{2})}{\gamma\alpha}.$ (4)
An example of the fit to the data can be seen in Figure 3. The constant offset $C_{2}$ is interpreted as being due to other loss mechanisms that scale proportionally with the superfluid density $\rho_{s}\propto f_{0}^{2}$. There are a number of candidates for this, which are discussed in the Supplementary Material [8].
Figure 3: Vortex Density Calibration Fit. (a) Plot of the experimental ratio $\Delta f/f_{0}^{2}$ as a function of temperature at 15 bar for the 636 nm device. This data has been fit to our model and plotted as a solid orange line. The dashed line displays the same curve if the offset is omitted in order to highlight the modification. (b) The vortex density is computed using the two methods shown in equation 4. Blue points are obtained by subtracting the fit offset $C_{2}$ from the data and dividing out pre-factors at each temperature. The orange line is computed directly from the slope $C_{1}$. The two methods are found to be consistent.
Discussion
Using this mutual friction model we extract a vortex density from the fit parameters of our Helmholtz resonance. We wish to compare this result to the predicted defect density of the Kibble-Zurek theory. It is worth pointing out that the confined dimension of the devices ($H=636-1067$ nm) is orders of magnitude smaller than the Kibble-Zurek length scale $\hat{\xi}=60-160$ $\mu$m. Because of this, we expect during the phase transition, the propagating domains should rapidly reach both walls. At this point, the domains cannot propagate further in the confined direction but should continue propagating in the two macroscopic directions. In this case the density of defects is determined by the Kibble-Zurek theory (see the Supplementary Material [8]) to be
$L=\frac{1}{\hat{\xi}^{2}}=\frac{1}{\xi_{0}^{2}}\left(\frac{\tau_{0}}{\tau_{Q}}\right)^{\frac{2\nu}{1+z\nu}},$ (5)
where $\xi_{0}$ is the zero temperature coherence length, $\tau_{0}$ is the relaxation time and $\tau_{Q}=T_{c}(dT/dt)^{-1}$ is a time constant characterizing the temperature ramp rate $dT/dt$. The exponents $\nu=1/2$ and $z=2$ are the spatial and dynamic critical exponents respectively. For the temperature ramp rate used in the majority of our experiments, 0.12 mK/hour, this corresponds to $L\sim 10^{7}-10^{8}$ m-2 over the range of pressures investigated in this experiment. Contrary to this, we measure vortex densities on the order of $L\sim 10^{10}-10^{11}$ m-2, as shown in Figure 4. Also notable is the clear dependence of vortex density on the confinement of the device; with the most confined channel ($636$ nm) having the highest density and the least confined ($1067$ nm) the lowest. Geometric effects do not enter into the original Kibble-Zurek theory, which should depend only on the temperature ramp rate and the pressure. Such high vortex densities (and therefore small domain sizes) would require extraordinarily fast quench rates (approximately 50-200 mK/s depending on the device) if they were dictated by the freeze-out time of the Kibble-Zurek theory.
In light of this discrepancy, we propose a modification to the Kibble-Zurek model. Consider a scenario where, $H<\hat{\xi}_{\textrm{KZM}}$, such that as the domain walls propagate outwards, they reach the wall of the device before reaching the freeze-out time. At this point, if a defect is created at the intersection of the domains, it may terminate at a wall rather than forming a loop. Unlike the loop, a wall-terminated line cannot shrink indefinitely to the point where it annihilates itself. The line is topologically protected such that the defect must remain even as the domains expand unless the ends can reconnect to form a loop. For this reason, we argue that the relevant lengthscale for the confined system is not $\hat{\xi}_{\textrm{KZM}}$, but an effective “truncation” length scale $\hat{\xi}=2H$ set by the wall separation (see Figure 4). By analogy to the Kibble-Zurek model, the density of defects would then be estimated as $L=H/\hat{\xi}^{3}=(8H^{2})^{-1}$. The truncation length prediction is found to be within an order of magnitude of the calibrated vortex density.
Figure 4: Comparison of the Kibble-Zurek theory and truncation model. (a) The prediction of the Kibble-Zurek theory, for a ramp rate of $dT/dt=0.12$ mK/hr, is plotted as a dashed black curve as a function of pressure. The superfluid coherence length, and therefore also $\hat{\xi}$, changes with pressure. This curve is compared to the experimental vortex densities — shown as a function of pressure for 636 nm (red), 805 nm (blue), and 1067 nm (green) devices — and found to differ by more than three orders of magnitude. The confinement-modified Zurek theory, on the other hand, well models the experimental vortex densities, as well as capturing the dependence on confinement, although it does not account for the pressure dependence found in the experiment. The x markers represent the fits to the raw data, and the circles are after correction for the artificial broadening of the Helmholtz resonances. (b) Cartoon diagram of a domain that propagates isotropically from an initial point to a distance $\hat{\xi}>2H$ (diameter), The walls of the domain become plane-like due to the existence of the confining walls. (c) A domain which has propagated a distance $\hat{\xi}=2H$. This is the point at which vortex lines joining the two walls first become possible.
Some degree of overestimation is expected as in the original Kibble-Zurek theory, since the function of the theory is to compute the average number of domain vertices, not all of which will have the necessary phase winding to become vortices. This is evidenced in numerical investigations of defect formation in phase transitions, which have found that the Kibble-Zurek prediction overestimates the defect density and that the true value is obtained by scaling $\hat{\xi}$ by a factor of 5-10 [36, 37, 38, 39, 40]. Such scaling has also been observed in 3He NMR experiments where, based on their measurements of the remanent vortex density, they estimate $\hat{\xi}$ should be scaled by 2.3 times the KZM value [28]. It is also interesting to note that these experiments have shown a suppression of defects compared to the KZ model, in contrast to the enhancement we see in the present experiments [41, 28]. Finally, we note that some vortices may annihilate at the boundaries, or due to vortex recombination during the phase transition [42]. We expect the latter mechanism to depend on the coherence length, which may explain the pressure dependence in the vortex density we observe. This is discussed in greater detail in the Supplementary Material [8].
Unlike the Kibble-Zurek model, where the vortex density is determined by a freeze-out time, our interpretation depends only on a spatial length-scale and is therefore ramp rate independent. After our initial measurements of the 636, 805, and 1067 nm devices, we later performed additional follow-up experiments using the 750 nm device to investigate any possible ramp rate dependence. This was achieved by varying the temperature ramp rate from 0.03-0.36 mK/hr while passing through the critical temperature. The maximum temperature ramp rates possible were limited both by the thermal time constant of the experimental cell, and the requirements that the nuclear demagnetization process used to cool our apparatus be adiabatic. Assuming the Kibble-Zurek scaling of $L\propto\sqrt{dT/dt}$, a factor of 10 increase in the ramp rate will increase the vortex density by $\sim$3 times, which should be observable in our experiment. Despite this, we see no evidence of a change in linewidth when the temperature is ramped at different rates through $T_{c}$. The only observed dependence on ramp rate was due to a systematic error introduced by our data averaging. Because many peaks were averaged together while the temperature was changing, the linewidth was artificially broadened. See the Supplementary Material for an in-depth discussion of the temperature ramp rate dependence [8]. In addition, we have performed measurements at fixed temperature, where we have cooled down without driving the device to ensure no turbulent generation of vortices, and find identical results, see Supplementary Material [8].
Conclusion
We have performed measurements of the dissipation of 3He fourth-sound resonances for several different confinements. We find that the temperature scaling of the dissipation is well-described by a vortex mutual friction model for a fixed vortex density. The lack of power dependence suggests that the density of vortices is set in upon phase transition rather than being extrinsically generated by flow (see discussion in Supplementary Materials [8]). We find that the density of vortices is independent of temperature ramp rate, but is a function of the channel thickness. A modified version of the Kibble-Zurek model is proposed in which a spatial length scale of the system sets the estimated defect density rather than the quench rate as is the case in the Kibble-Zurek model. This interpretation is found to be within an order of magnitude of the observations, comparable to the agreement seen in tests of the Kibble-Zurek mechanism, and the scaling with confinement is in reasonable agreement. We believe this result demonstrates the non-trivial dependence of defect density on the boundaries of the system. This may be relevant to a variety of systems where the Kibble-Zurek length scale is small compared to a single dimension of the system.